OpenStudy (anonymous):

Using complete sentences, explain the two ways to simplify the square root of 320 and provide the simplified form. Need someone to check my answer, I will post it below...

OpenStudy (anonymous):

Find the largest square number that goes in to 320, i.e. 64 and rewrite 320 as a multiplication 64 and another number. The square root of 320 is the same as the multiplication of the square roots of both numbers individually. Since the square root of a square number will always be a whole number, it can be simplified and multiplied by the square root of the other number. √(320) = √64*√5 = 8√5 You could also take the longer route (take any square you can find): √(320) = √4*√16*√) = 2*4√5 = 8√5 is this right and does it make sense?

OpenStudy (anonymous):

looks good.

OpenStudy (anonymous):

yay! thanks :)

OpenStudy (anonymous):

and to you, thanks, because you're the one that gave me my new smartscore!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

yw too...